Your child is planning to attend summer camp for 3 months, starting 6 months from now.
The cost of the camp is $4,000 per month, each month for the 3 months she will attend.
If your investments earn 5% APR (compounded monthly) how much must you invest each month
starting next month, for 3 months, such that your investment would just cover the cost of the camp?
Answer #1

Total cost for attend summer camp = $12000

invest each month

pv = $12000/(1+.005/4)3/12 = $12000*.9515242752 = 11418

per month = 11418/3 = 3806 invest